# FMANet: A Novel Dual-Phase Optical Flow Approach with Fusion Motion Attention Network for Robust Micro-expression Recognition

### v2

@@ Abstract
Facial micro-expressions, characterized by their subtle and brief nature, are valuable indicators of genuine emotions. Despite their significance in psychology, security, and behavioral analysis, micro-expression recognition remains challenging due to the difficulty of capturing subtle facial movements. Optical flow has been widely employed as an input modality for this task due to its effectiveness. However, most existing methods compute optical flow only between the onset and apex frames, thereby overlooking essential motion information in the apex-to-offset phase. To address this limitation, we first introduce a comprehensive motion representation, termed Magnitude-Modulated Combined Optical Flow (MM-COF), which integrates motion dynamics from both micro-expression phases into a unified descriptor suitable for direct use in recognition networks. Building upon this principle, we then propose FMANet, a novel end-to-end neural network architecture that internalizes the dual-phase analysis and magnitude modulation into learnable modules. This allows the network to adaptively fuse motion cues and focus on salient facial regions for classification. Experimental evaluations on the MMEW, SMIC, CASME-II, and SAMM datasets, widely recognized as standard benchmarks, demonstrate that our proposed MM-COF representation and FMANet outperforms existing methods, underscoring the potential of a learnable, dual-phase framework in advancing micro-expression recognition.
@@ II Magnitude-Modulated Combined Optical Flow Representation
Micro-expressions unfold through asymmetric temporal phases: the onset–apex phase captures the emergence and peak intensity, while the apex–offset phase reflects relaxation and transition. Assuming symmetry between these phases oversimplifies the dynamics and risks discarding informative offset cues. This challenge is exacerbated by the fact that apex frames are rarely centered between onset and offset in existing MER datasets.
@@ III Fusion Motion Attention Network
III-1 Phase-Aware Consensus Fusion Block
@@
III-2 Soft Motion Attention
@@ IV-A Dataset
The datasets used in this study are CASME-II [29], SAMM [30], SMIC [31] and MMEW [32], both widely recognized as standard benchmarks in MER research due to their high-quality samples.
@@ IV-C Evaluation Metrics
Given the inherent class imbalance in micro-expression datasets, multiple metrics are employed to provide a comprehensive assessment of model performance.
@@ IV-D1 Visual representation results
The results indicate that both the FFB and SMAB blocks successfully fulfill their intended purpose. The FFB effectively combines homogeneous regions from the two motion phases, and the SMAB concentrates on the critical regions within the FFB’s feature representation. Specifically, compared to the original optical flow representations, the motion features generated by the FFB are more effective at reducing noise and focusing on significant areas. For example, in the ”Surprise” and ”Happiness” expressions, the movements are subtle and tend to produce considerable noise in the optical flow representation. However, our method successfully distills the core motion information by unifying the two motion phases, and the SMAB effectively targets these key regions.
Furthermore, the results demonstrate that with the representations generated by the FFB and SMAB, the corresponding Action Units can be observed and identified much more easily and accurately than with the original optical flow representation.
IV-D2 Comparison with State-of-the-arts
@@ IV-E3 Analysis of Magnitude-modulated Combined Optical Flow (MM-COF)
Input modalities: Finally, Table XI compares MM-COF with other representations. It is observed that magnitude-modulated representations (MM-COF) consistently outperform plain optical flow and apex frame inputs. Among the backbones, SCNN delivers higher and more stable performance than VGG19 across both datasets, confirming that a compact, task-specific architecture is better suited for capturing subtle micro-expression motions than a generic deep model. Notably, MM-COF with SCNN achieves the best single-stream results, reaching 70.08% on CASME-II and 63.70% on SAMM.
TABLE XI: Evaluation of different input modalities in the single-stream setting using VGG19 and SCNN on the CASME-II and SAMM datasets. Results are reported in Accuracy (%).
Input modality CASME-II SAMM
VGG19 SCNN (ours) VGG19 SCNN (ours)
Apex frame 59.59 63.27 55.15 54.41
Optical flow 54.10 63.52 58.15 54.41
COF 56.97 65.98 58.52 59.26
$M$-COF 63.93 70.08 60.74 63.70

### v3

@@ Abstract
Facial micro-expressions, characterized by their subtle and brief nature, are valuable indicators of genuine emotions. Despite their significance in psychology, security, and behavioral analysis, micro-expression recognition remains challenging due to the difficulty of capturing subtle facial movements. Optical flow has been widely employed as an input modality for this task due to its effectiveness. However, most existing methods compute optical flow only between the onset and apex frames, thereby overlooking essential motion information in the apex-to-offset phase. To address this limitation, we first introduce a comprehensive motion representation, termed Magnitude-Modulated Combined Optical Flow (MM-COF), which integrates motion dynamics from both micro-expression phases into a unified descriptor suitable for direct use in recognition networks. Building upon this principle, we then propose FMANet, a novel end-to-end neural network architecture that internalizes the dual-phase analysis and magnitude modulation into learnable modules. This allows the network to adaptively fuse motion cues and focus on salient facial regions for classification. Experimental evaluations on the CASME-II, SAMM and MMEW datasets, widely recognized as standard benchmarks, demonstrate that our proposed MM-COF representation and FMANet outperforms existing methods, underscoring the potential of a learnable, dual-phase framework in advancing micro-expression recognition.
@@ II Magnitude-Modulated Combined Optical Flow Representation
Micro-expressions unfold across two distinct temporal phases: an onset-to-apex (formation) phase and an apex-to-offset (relaxation) phase. While the complete motion in these phases is not perfectly identical, exhibiting asymmetry due to non-key motions, subtle variations in muscle relaxation, or temporal shifts, we posit that the core muscle activations defining the genuine emotion are fundamentally symmetric. Therefore, our fusion of the two stages is motivated by the goal of filtering out the asymmetric non-key motions while retaining and reinforcing only the symmetric key motions. This approach allows us to distill a more robust and purified representation of the true expression, a challenge exacerbated by the fact that apex frames are rarely centered in existing MER datasets.
@@ II-A Optical Flow Calculation
The onset, apex, and offset frames serve as critical anchor points that capture the entire life-cycle of a micro-expression: its formation, peak intensity, and subsequent relaxation. Therefore the optical flow from the two phases will be calculated by frame onset and apex, frame apex and offset.
@@ III Fusion Motion Attention Network
III-A Phase-Aware Consensus Fusion Block
@@
III-B Soft Motion Attention
@@ IV-A Dataset
The datasets used in this study are CASME-II [29], SAMM [30] and MMEW [31], both widely recognized as standard benchmarks in MER research due to their high-quality samples.
@@ IV-C Evaluation Metrics
These metrics are employed to provide a comprehensive assessment of model performance:
@@ IV-D1 Visual representation results
The results indicate that both the FFB and SMAB blocks successfully fulfill their intended purpose. The FFB effectively combines homogeneous regions from the two motion phases, and the SMAB concentrates on the critical regions within the FFB’s feature representation. Specifically, compared to the original optical flow representations, the motion features generated by the FFB are more effective at reducing noise and focusing on significant areas. For example, in the ”Surprise” and ”Happiness” expressions, the movements are subtle and tend to produce considerable noise in the optical flow representation. However, our method successfully distills the core motion information by unifying the two motion phases, and the SMAB effectively targets these key regions. Furthermore, the results demonstrate that with the representations generated by the FFB and SMAB, the corresponding Action Units can be observed and identified much more easily and accurately than with the original optical flow representation.
@@ IV-E3 Analysis of Magnitude-modulated Combined Optical Flow (MM-COF)
Input modalities: Table XI further compares MM-COF against conventional representations using three backbone networks: VGG19, ConvNeXtV2, and the proposed SCNN. VGG19 is included as a widely adopted baseline CNN with a deep yet generic architecture, while ConvNeXtV2 represents a more recent convolutional design optimized for large-scale visual recognition. Despite their capacities, both models are not specifically tailored for capturing subtle and low-amplitude facial motions. In contrast, SCNN demonstrates superior and more consistent performance across both datasets, highlighting the advantage of a compact, task-focused architecture for micro-expression analysis. Moreover, MM-COF paired with SCNN yields the highest single-stream results, achieving 70.08% on CASME-II and 63.70% on SAMM, confirming the effectiveness of motion modulation in enhancing discriminative power.
TABLE XI: Evaluation of different input modalities using VGG19, ConvNeXtV2, and SCNN on CASME-II and SAMM datasets (Accuracy %).
Input Modality Model CASME-II SAMM
Apex frame VGG19 59.59 55.15
ConvNeXtV2 60.66 57.04
SCNN (ours) 63.27 54.41
Optical flow VGG19 54.10 58.15
ConvNeXtV2 59.84 56.30
SCNN (ours) 63.52 54.41
COF VGG19 56.97 58.52
ConvNeXtV2 61.89 57.78
SCNN (ours) 65.98 59.26
$M$-COF VGG19 63.93 60.74
ConvNeXtV2 64.75 60.00
SCNN (ours) 70.08 63.70